OpenStudy (ccastudentgrade10):

PLEASE HELP ASAP write the equation in slope intercept form -9x + 10y = -9

OpenStudy (anonymous):

9x-9 =10y 0.9x - 0.9 =y

OpenStudy (denonakavro):

y=(9/10)x - (9/10)

OpenStudy (ccastudentgrade10):

thanks guys, what's the y-intercept then? @lall @denonakavro